### v1

# Fast algorithms to improve fair information access in networks

Abstract
When information spreads across a network via pairwise sharing, large disparities in information access can arise from the network’s structural heterogeneity. Algorithms to improve the fairness of information access seek to maximize the minimum access of a node to information by sequentially selecting new nodes to seed with the spreading information. However, existing algorithms are computationally expensive. Here, we develop and evaluate a set of 10 new scalable algorithms to improve information access in social networks; in order to compare them to the existing state-of-the-art, we introduce both a new performance metric and a new benchmark corpus of networks. Additionally, we investigate the degree to which algorithm performance on minimizing information access gaps can be predicted ahead of time from features of a network’s structure. We find that while no algorithm is strictly superior to all others across networks, our new scalable algorithms are competitive with the state-of-the-art and orders of magnitude faster. We introduce a meta-learner approach that learns which of the fast algorithms is best for a specific network and is on average only 20% less effective than the state-of-the-art performance on held-out data, while about 75-130 times faster. Furthermore, on about 20% of networks the meta-learner’s performance exceeds the state-of-the-art.
I Introduction
Fairness in machine learning is a well studied area, and includes notions of fairness based both on individual characteristics and on group demographics feldman2015certifying ; hardt2016equality ; dwork2012fairness (for surveys, see Refs. caton2024fairness ; pessach2022review ). However, fairness in the context of social networks is comparatively less studied. Individual Fish_2019 and group stoica2019fairness notions of fairness have been introduced that assess individuals’ access to information that spreads across a social network. In this work, we introduce fast algorithms to achieve individual fairness in networks, as defined by Ref. Fish_2019 .
In social networks, access to information is unequally distributed. This inequality is a natural consequence of heterogeneous network connectivity and position—highly-connected or centrally located individuals have more opportunities to receive and spread information cheng2019 ; PICARD201877 ; toth2021 , while peripheral individuals with few connections participate less often in information exchanges PICARD201877 ; Fish_2019 . For example, in a pandemic, access to crucial resources—such as money, food and healthcare—is more difficult for socially disadvantaged groups, in part due to their more limited connectedness in social networks GAROON20081133 . Similarly, connectedness shapes employment opportunities in professional social networks such as LinkedIn: well-connected job seekers are likely to fill lucrative openings sooner than others weaktiesinlinkedin2022 .
The quality of a node’s information access can be formalized mathematically. Given a graph $G=(V,E)$, let $S\subset V$ be the seed set, where $x\in S$ represents an individual who is seeded with information and propagates it across the network via a word-of-mouth process. For a node $i\in V$, we say that the probability $\pi_{i}$ that node $i$ receives the information measures the information access of the individual $i$ under $S$. The notion of individual fairness on networks we will consider Fish_2019 seeks to maximize the minimum such information access $\pi_{i}$ over all individuals $i\in V$, with the aim of focusing interventions on the most disadvantaged with respect to information access in the network.
Figure 1: Algorithm runtime to select 10 new seeds vs. network size for algorithms in Ref. Fish_2019 , averaged over 10 runs on an introduced large set of networks (see Section III.1). Algorithms requiring a Monte Carlo simulation (ProbEst) to select seeds are denoted by a *.
Selecting $k$ nodes at which to seed information to maximize $\min_{i\in V}\pi_{i}$ is NP-hard, and hence past research has proposed heuristics Fish_2019 . However, the heuristics’ performance has been evaluated on a structurally limited range of networks Fish_2019 , and it remains unclear how algorithm performance depends on the structure of the network. Additionally, the algorithms rely on computing $\pi_{i}$ based on the independent cascade model for information propagation, but it is #P-hard to do so exactly provan1983complexity . Therefore, most research has relied on Monte Carlo simulation (here denoted ProbEst; see Section IV.1) to estimate $\pi_{i}$ for each node in the network Fish_2019 . However, ProbEst has high computational cost, making it impractical on very large networks (see Fig. 1). In addition, algorithms that rely on ProbEst can only be evaluated after choosing a value for the parameter $\alpha$, i.e., the independent cascade model’s probability of successful transmission across any given edge kempe_influence_maximization_2003 . The choice of $\alpha$ is not trivial, as the structure of the network influences how well information spreads under a given $\alpha$, and there exists no standard metric for comparing algorithm performance on a given network, or comparing across networks. Finally, it remains unknown whether new, more scalable algorithms can achieve comparable performance to algorithms based on ProbEst, and it is unknown whether one algorithm always generates the best seed sets regardless of the network, and if not, how we might choose the best algorithm given only characteristics of a network.
Progress on these questions would advance our understanding of how to, given a budget constraint and a target network, devise an optimal intervention technique that serves to improve information access of the disadvantaged individuals. This knowledge would be instrumental in designing effective public health campaigns, outreach strategies, etc., that aim to improve fairness in social networks by reaching even the most socially isolated individuals alexander_india_healthcare_2022 . This goal is particularly important because conventional influence maximization techniques have been shown to be ineffective in improving information access of the disadvantaged nodes Fish_2019 .
I.1 Contributions
In our work, we introduce ten new fast algorithms to improve fairness in information access; specifically to maximize the minimum $\pi_{i}$ given a budget of $k$ nodes at which to seed information. We introduce a comprehensive cross-domain network corpus and introduce the concept of spreadability, which we define as the ability of a network to spread information under a given $\alpha$. We subsequently pick $\alpha$ values individually for each network based on a target spreadability. We introduce evaluation mechanisms to assess the quality of fair information access algorithms including a new metric $\beta$, the average marginal increase in access for each additional seed selected. Using the introduced corpus and evaluation mechanisms, we assess the new fast algorithms in comparison to the state-of-the-art. We provide the network corpus and code for this study at https://github.com/rhythmthief/FairnessNetworks. Finally, we introduce a meta-learner approach that chooses the best fast algorithm based on network structure features alone, and is, on average, only 20% less effective than the state-of-the-art while being 75-130 times faster. Furthermore, on about 20% of networks the meta-learner’s performance exceeds the state-of-the-art.
II Related Work
Considerations about the access of individuals to resources in a network build on structural concerns about social networks pioneered by Ref. granovetter1973 . When considering how resources are shared across a network, two popular approaches to spreading information are broadcasting and word-of-mouth NBERw24679 ; kempe_influence_maximization_2003 ; watts_influentials_2007 . Broadcasting is a one-to-many approach, where a single source sends information to all nodes in the network. Word-of-mouth is a many-to-many approach, where information is spread through the network by individuals sharing it with their neighbors NBERw24679 . The latter approach leverages the structure of the network to spread information, and real-world data suggests it can be more effective than broadcasting in certain settings NBERw24679 .
Numerous models of information propagation through networks exist, including the independent cascade model, generalized independent cascade model, and linear threshold model kempe_influence_maximization_2003 ; granovetter1978 ; goldenberg2001 . A popular choice in prior work, also adopted by us, is the independent cascade model, which encapsulates the fundamental dynamics of information spread in networks kempe_influence_maximization_2003 ; Fish_2019 . It operates stochastically, where each edge may facilitate a single attempt to transmit information from an active node to its inactive neighbor, with some probability $\alpha$ of success kempe_influence_maximization_2003 . The independent cascade model’s simplicity provides an effective framework for studying information dissemination across networks.
Under a word-of-mouth approach, information access in a network can be improved via selective seeding, i.e., by choosing an individual or set of individuals to be the first to receive the information NBERw24679 . This idea also underlies the field of influence maximization, which aims to identify the best seed set in order to maximize the spread of information through the network kempe_influence_maximization_2003 . A variety of algorithms have been proposed to solve the influence maximization problem kempe_influence_maximization_2003 ; SINGH20227570 ; Zareie2023 . In contrast to our work, the traditional goal has been to maximize the expected number of activated nodes by selecting a small set of initially activated individuals to seed an information cascade kempe_influence_maximization_2003 . This approach does not account for the fairness of the spread among individuals in the network.
Recently, questions of fairness of information access of individuals and demographic groups in a social network have come to the fore Fish_2019 ; stoica2019fairness ; stoica2020seeding ; ali2021fairness ; mehrotra2022revisiting ; tsang2019group-fairness ; stoica2018algorithmic . Much work focuses on the probability that an individual receives information that spreads through a social network. This work includes studies of seeding information at nodes to improve individual access to that information Fish_2019 , or group access tsang2019group-fairness ; stoica2020seeding ; or interventions to add edges bashardoust2023reducing ; becker2023improving ; bhaskara2024optimizing under varying notions of fairness. Other work considers notions of group fairness based on the network structure jalali2023fairness ; mehrotra2022revisiting ; bashardoust2020information . Ref. saxena2024fairsna provides a recent survey of such work.
Figure 2: Average degree of a network as a function of network size (number of nodes) for the corpus of 174 networks from 6 distinct domains used in our study.
III Network Corpus and Evaluation Methods
III.1 Network Corpus
In order to investigate the effects of network structure on algorithm performance, we construct a corpus of $174$ networks from six domains: biological, social, economic, technological, transportation, and informational. We include non-social networks in our study in order to more fully characterize the behavior of algorithms on structurally diverse real-world networks. When relevant, we report our results by domain, so that social and non-social networks can be contrasted.
The networks in the corpus have at least 500 nodes, and are unipartite, that is, they only have one type of node. We simplified each network by ignoring edge weights, edge direction and self-loops, and took the largest connected component of each network, which itself must contain at least 500 nodes. We also ensured that no domain accounts for more than 25% of all networks in the corpus. Data for the corpus was obtained from  Ghasemian_2019 ; ICON ; Netzschleuder . An overview of the corpus is presented in Figure 2, with summary statistics in Table S1.
III.2 Spreadability
In our work, we leverage the independent cascade model to study the spread of information in networks and estimate $\pi_{i}$. Given a network $G$ and a set of activated nodes $Q\subseteq V$, we grow a forest on $G$ under the independent cascade model by flipping a coin for each edge $e_{ij}$, where $i\in V\backslash Q$, $j\in Q$, exactly once, so that $i$ is added to $Q$ on a successful flip. Each edge $e_{ij}$ is considered as a transmission path at most once, and we define the probability of successful transmission to be $\alpha$. We note that mathematically, it is equivalent to conceptualize the independent cascade as follows: delete every edge in the graph independently with probability $(1-\alpha)$; the set $Q$ of activated nodes is exactly the set of nodes that remain reachable from the seed set $S$. As noted earlier, computing the exact probability of activation for a node $i$, denoted $\pi_{i}$, is #P-hard. As such, we adopt the standard Monte Carlo simulation approach, ProbEst Fish_2019 , but return later to consider practical consequences of this choice.
Figure 3: Spreadability on a network is quantified by the average fraction of a network’s nodes $\langle|T|\rangle/n$ in a tree $T$ grown through an independent cascade from a random initial seed for a given $\alpha$. We define ‘low’, ‘medium’, and ‘high’ spreadability as the $\alpha$ that activates, on average, $20\%$, $50\%$, and $80\%$ of the network, respectively.
Briefly, using ProbEst, given transmission probability $\alpha$, number of simulation rounds $R$, and a seed set $S$, we estimate $\pi_{i}$ for every $i\in V$ using $R$ independent cascades originating from $S$ Fish_2019 . The worst-case time complexity of ProbEst is $\mathcal{O}(R(|S|+2m))$ Fish_2019 , where $m$ is the number of edges in $G$. In practice, the computational cost of ProbEst increases both as the seed set grows and as $\alpha$ increases, because both changes tend to increase the size of the induced information cascades. Past work used $R=1000$ as a balance between statistical accuracy for $\pi_{i}$ and computational cost, which we incorporate in our work as well Fish_2019 . The output of ProbEst is used here both as part of some evaluated algorithms and to evaluate their performance.
Prior work evaluated algorithm performance using transmission probabilities $\alpha\in\{0.3,0.4,0.5\}$ Fish_2019 . However, the resulting information cascades, and hence the associated access probabilities, are not a simple function of $\alpha$; they instead also depend on the network’s structure. For example, denser, more connected networks contain many more paths by which information can spread than do sparse networks. Hence the same $\alpha$ will tend to produce larger cascades on the former, and smaller cascades on the latter. To control for these structure-induced differences in information cascades, we introduce the concept of spreadability, which jointly accounts for the impact of network structure and transmission rate $\alpha$ on the sizes of information cascades.
Figure 4: Minimum access probability $\pi_{\min}$ vs. seed set size $k$, with a best fit line (Myopic $\hat{\beta}=0.039$), averaged over 20 runs, evaluated on a large economic network ($n=2113$ nodes, $m=57927$ edges), with $\alpha=0.4$ and a budget of $k=10$ seeds, plus one random initial seed.
We can construct a fine-grained spreadability function that relates a particular choice of $\alpha$ to the fraction of a network activated under the independent cascade model from a uniformly random initial seed. For a given $\alpha$, the spreadability $f(\alpha)$ on a particular network is computed by averaging over $R$ trials the fraction of nodes that are activated from a uniformly random initial seed. This calculation produces a monotonically increasing curve, as seen in Figure 3.
We find that computing spreadability for each $\alpha\in\{0.01,0.02,...0.99\}$ provides ample resolution to choose $\alpha$ close to target spreadabilities of $0.2,0.5,0.8$ (which we refer to as “low,” “medium,” and “high” spreadabilities respectively). We set $R=1000$, as the spreadability curve tends to stabilize near this value and we get diminishing returns for larger $R$.
III.3 New Metric for Algorithm Evaluation
We propose a metric for evaluating the relative performance of algorithms, $\beta$, which corresponds to the average slope of the lines of best fit for ProbEst-based evaluations after selecting the first $10$ seeds, as seen in Figure 4. This new metric is useful because it shows the marginal improvement in $\pi_{\rm min}$ that we can expect when asking an algorithm to draw a new seed, which allows for a direct comparison of the performance of the algorithms. We use this metric to evaluate the performance of our algorithms on the network corpus.
IV Algorithms for Fair Information Access
IV.1 Algorithms from Prior Work
Four previously introduced heuristics for choosing seed nodes to maximize $\pi_{\rm min}$ are Greedy, Myopic, Naive Myopic and Gonzalez, along with Random as a baseline for comparison Fish_2019 . Random selects $k$ seeds by uniformly sampling nodes from the network. Given a partial seed set, Gonzalez selects as the next seed the node that is the furthest (in the shortest-path metric) from all nodes in the current set Fish_2019 . The Greedy algorithm iteratively selects a new seed by choosing the node with the highest marginal gain relative to the current seed set according to ProbEst. Due to its immensely high computational cost, Greedy is not practical for most networks Fish_2019 , and is not used in our study.
In contrast, Myopic uses ProbEst with the current seed set, and selects as the next seed the node with the lowest $\pi_{i}$. Naive Myopic is similar to Myopic, but only runs ProbEst once, at initialization, and then selects as the seed set the nodes with the lowest $\pi_{i}$ values. In past work, on a small set of networks, Myopic was found to perform best. However, because Myopic depends on ProbEst, it is computationally expensive, which limits its applicability to large networks. In practice, Myopic is always the second slowest algorithm, after Greedy Fish_2019 .
Figure 5: Mean performance of the intervention algorithms on each domain in the corpus under medium spreadability. Each algorithm’s performance is averaged over a given domain, 20 runs per network.
Algorithm Initialization
Myopic, Naive Myopic, and Gonzalez all start with an initial seed. Past work chose this initial seed to be the highest degree node. Our initial experiments indicate that this choice confers a substantial advantage to these methods (Fig. S1), and that some of the previous positive results are thus attributable to this initial seed choice rather than to the algorithms’ subsequent choices. To mitigate this bias we instead initialize all heuristics with a seed set composed of a single uniformly randomly selected node. Moreover, for each evaluation round on a particular network, we initialize all algorithms to use the same random seed, which controls for the effects of different initial seeds. This initial seed choice is not counted against the budget $k$ (Fig. 4).
IV.2 New Fast Algorithms
In addition to the four heuristics from prior work, here we introduce 10 new heuristics to maximize the minimum $\pi_{i}$ on a budget. These heuristics are designed to be computationally lightweight, scaling to far larger networks, while also matching or exceeding the performance of Myopic on the corpus. The new algorithms can be grouped into three families: BFS-based, PPR-based and Topology-based.
BFS-based: Myopic BFS and Naive Myopic BFS
In the two BFS-based heuristics, Myopic BFS and Naive Myopic BFS, we swap the ProbEst component of Myopic and Naive Myopic with a simple breadth-first search to estimate $\pi_{i}$. The breadth-first search component is initialized with a random seed $k_{0}$, and transmission probability $\alpha$. It proceeds to “peel” the network, starting at $k_{0}$, in breadth-first fashion, estimating $\pi_{i}$ for each node as the probability that $i$ receives a transmission from $k_{0}$ through any nodes it connects to in the previous BFS layer, as well as through nodes it is connected to in its own layer. All subsequent iterations update existing $\pi_{i}$ estimates during the breadth-first traversal from new candidate seeds.
While this approach does not exactly measure $\pi_{i}$, it is much faster than ProbEst, taking $\mathcal{O}(n\cdot\langle k\rangle)$ per iteration, because for most networks, including those in our corpus, the mean degree $\langle k\rangle\ll 1000$ (Fig. 2; here, $k$ denotes the degree of a node). The key design principle of this algorithm is to capture complex network structures that Gonzalez could not account for. In Figure 4, we find that Myopic BFS almost matches Myopic’s performance in situations where Gonzalez lags behind.
PPR-based: Myopic PPR and Naive Myopic PPR
In the PPR-based heuristics, Myopic PPR and Naive Myopic PPR, we use Personalized PageRank (PPR) to estimate $\pi_{i}$ instead of ProbEst or BFS. Personalized PageRank (implemented with networkx SciPyProceedings_11 ) performs a random walk that probabilistically restarts from nodes in the seed set PPR1998 . The ranking produced by PPR is not a direct estimate of $\pi_{i}$; we treat the PPR values as being correlated, such that a lower PPR score is a proxy for a lower $\pi_{i}$ value.
Topology-based: LeastCentral and MinDegree variants
The topology-based heuristics are based on the intuition that nodes with low values of $\pi_{i}$ have distinctive structural positions or patterns of connectivity. LeastCentral selects the non-seed node $i$ with the lowest closeness centrality $c_{i}$ as the next seed. Similarly, LeastCentral_n selects the lowest centrality node $i$’s highest degree neighbor as the next seed. Here, the closeness centrality of a node $i$ is the inverse of the average shortest path length from $i$ to all other reachable nodes in the same connected component FREEMAN1978215 ; newman18 . Lower closeness centrality implies the node is less reachable by the rest of the network, and therefore is expected to have lower $\pi_{i}$.
The four remaining topology-based heuristics exploit a network’s degree structure to make decisions, based on the observation that in practice Myopic tends to select nodes with low degree as seeds. The first two heuristics take a non-seed node with the lowest degree, breaking ties by choosing the node with the lowest harmonic centrality boldi2013axioms ; newman18 among same-degree nodes. In the case of MinDegree_hc, it chooses that node itself as the next seed, while MinDegree_hcn chooses the highest-degree neighbor of that node. The two remaining variations replace harmonic centrality in the aforementioned logic with the neighbor degree, i.e. breaking ties by choosing the node with the highest neighbor degree (sum of degrees across neighbors). These heuristics are called MinDegree_nd and MinDegree_ndn, respectively.
Figure 6: Performance of intervention algorithms on the network corpus, relative to Myopic and sorted in ascending order by network size within each domain. Under medium spreadability, $24\%$ of networks have no algorithm better than or within $80\%$ of Myopic’s performance. “Equivalent” is defined as within one standard error of $\beta$ for Myopic; typically about 0.001.
V Experimental Results
V.1 Performance of Algorithms on the Corpus
We evaluate and compare the performance of 14 algorithms in total (10 new algorithms and 4 from prior work), applied to all $174$ networks in the corpus. We are interested in algorithms that operate well on a tight budget and so let $k=10$ seed nodes. For each of the three spreadability levels (low, medium, high), we produce a $10\times 14\times 174$ matrix, where each entry is the $\beta$ performance of an algorithm after adding $k_{n}\in\{1,\dots,10\}$ seeds in a network, averaged over $20$ runs. We focus on the medium spreadability here, and include results for low and high spreadability in the appendix. Figure 5 displays the mean performance of each algorithm on networks in a given domain under medium spreadability (see Fig. S2 for low and high spreadability results). Across domains, Myopic produces the best average performance.
To further examine the results, we sort within each domain by network size in ascending order, and then score algorithms as better than, “equivalent” to (within one std. err.), within 80% of, or worse than Myopic (see Fig. 6; low and high spreadability results in Fig. S3). In this way, we can assess whether the better average performance of Myopic applies to individual networks compared to other algorithms. This experiment reveals that Myopic is not universally the best algorithm for all networks in any spreadability setting; on many individual networks other algorithms perform equivalently or better. In the medium spreadability setting, for only $24\%$ of networks is there no algorithm that performs at least $80\%$ as well as Myopic (Fig. 6).
This variability in performance across networks suggests that network structure plays a critical role in governing the relative performance of different algorithms. Figure 7 plots the instances for which each algorithm was the best performing algorithm on a network against that network’s mean degree $\langle k\rangle$. Myopic tends to perform better on networks with lower average degree, although it does also perform well for many networks with larger mean degree. In practice, we find that the final seed set produced by Myopic is often composed primarily of low-degree nodes located in a network’s periphery far from one another, and these nodes may be too far removed from other disadvantaged nodes to meaningfully improve their $\pi_{i}$ values.
A second takeaway is that a few specific algorithms tend to perform better than Myopic in certain settings (Fig. 7), specifically MinDegree_hcn and Gonzalez. The performance of MinDegree_hcn in particular tends to improve over Myopic with increasing average degree, while Gonzalez does best in networks with very low mean degrees. Furthermore, we note that in Figure S2A, on average MinDegree_hcn outperforms Myopic in the economic domain, and from Table S1, we see that the economic domain has the highest mean degrees. The MinDegree_hcn algorithm selects as seeds the highest-degree neighbors of low-degree nodes. As result, new information cascades seeded at these nodes will tend to spread quickly to a number of disadvantaged nodes in the network.
Figure 7: Best-performing algorithm vs. mean degree of the network (medium spreadability), for all networks. Counts on the right show total circles per line, i.e., the number of times an algorithm was the best over the whole corpus.
V.2 Algorithm Runtime
We evaluate all algorithm runtimes on the corpus, selecting $k\!=\!10$ new seeds, averaged over $10$ runs. For fair comparison, we run all algorithms on a single core of an AMD Ryzen 5900X, overclocked to $5.00$Ghz, with 32GB RAM, and measure the runtime in milliseconds. Two major performance bottlenecks are ProbEst, used by Myopic and Naive Myopic, and an All-Pairs-Shortest-Paths (APSP) computation, used by Gonzalez, LeastCentral, LeastCentral_n, MinDegree_hc, and MinDegree_hcn. Both ProbEst and APSP have efficient parallel implementations, but we restrict them to a single core to ensure fair comparison with other algorithms.
All of the new algorithms are substantially faster than Myopic and Naive Myopic (Fig. 8). Because they only require sorting two lists, MinDegree_ndn and MinDegree_nd are the most efficient, improving running times over ProbEst-based algorithms by a factor of 1000-10000x, depending on the size of the network (Fig. 8). BFS-based algorithms are marginally slower than these fastest algorithms, and algorithms that use an All-Pairs-Shortest-Paths (APSP) calculation fall between the ProbEst and BFS algorithm groups. As we would expect from the asymptotics, BFS algorithms tend to scale more slowly (in terms of runtime) with network size than do ProbEst algorithms, while APSP algorithms scale more quickly (Fig. 8).
The low upfront cost of APSP-based algorithms makes them far faster than ProbEst-based algorithms in many practical settings, being 10-100x faster on networks with less than $n=10^{6}$. However, the asymptotic cost of APSP grows super-linearly, implying that for sufficiently large networks, ProbEst will be faster. For our corpus, we estimate the crossover point when Myopic becomes more efficient than MinDegree_hc to occur between $1,262,000$ and $1,515,000$ nodes (95% CI, from 1000 bootstraps). However, approximation algorithms for APSP could potentially extend their practical efficiency much further.
Figure 8: Runtime of old and new algorithms across the network corpus with $k=10$ seeds averaged over 10 runs, showing a substantial advantage in running time for the new algorithms. Note: algorithms with * require ProbEst.
V.3 A Fast Meta-Learning Algorithm
We can exploit the variability in algorithm performance, and the fact that even among non-Myopic algorithms no alternative is superior on all networks, by introducing a meta-learner algorithm that combines multiple scalable heuristics to approximate the state-of-the-art performance of the Myopic algorithm. We compare this algorithm to a fast ensemble algorithm that uses an oracle to make perfect predictions about which scalable algorithm is best to apply on a particular test network, thus providing an upper bound on the meta-learner’s possible performance.
The meta-learner algorithm leverages the scalability of non-ProbEst algorithms while retaining good overall performance under all spreadability regimes. The task is as follows: given a particular network $G$ and knowledge of the information’s spreadability (low, medium, or high), select the scalable algorithm with the best marginal benefit $\beta$ for improving information access.
As shown previously, many of the heuristics do not perform well, and so we begin by narrowing the set of available algorithms. For each of the three spreadability settings, we select the set $\mathcal{X}$ of five algorithms (excluding Myopic and Naive Myopic) that maximize the number of networks for which at least one among the set performs at least $80\%$ as well as Myopic. This produces sets
[…]
For the meta-learner algorithm, we then learn a 5-way random forest classifier to predict the best algorithm in $\mathcal{X}$ to apply to a given network, using nine of the network’s topological features as the feature set (Fig. S6). We train and evaluate the meta-learner approach using an 80-20 train-test split among networks in the corpus, with meta-learner algorithm selection and model training both performed on the training set, and we report the mean performance over the test set. The meta-learner’s runtime is the runtime of the trained model and the single algorithm it selects.
In the fast ensemble algorithm, for a given network in the corpus, the oracle runs all algorithms in $\mathcal{X}$ and evaluates the performance of each using ProbEst to calculate their respective $\beta$s, and then returns the single algorithm with the highest $\beta$ for that network. In this way, the oracle acts like an optimal classifier over $\mathcal{X}$ (cf. the meta-learner algorithm). The runtime of the fast ensemble algorithm is simply that of the single algorithm it selects.
Figure 9: Performance difference vs. speedup for the meta-learner algorithm under medium spreadability, with marginal histograms, averaged over $1000$ runs. Extreme outliers have been removed for visualization purposes. Average performance difference relative to Myopic is $-20.11\%\pm 29.34$ (mean $\pm$ stddev), for an average speedup factor of $76.28\pm 64.07$. For 34 of the networks (19.8%) the meta-learner strictly outperforms Myopic.
Compared to Myopic, the meta-learner is dramatically more efficient, with an average runtime that is $76.26\pm 64.07$ times faster under medium spreadability (Fig. 9) and $133.35\pm 79.32$ times faster under high spreadability (Fig. S8). Improvement in scalability comes with a modest cost to performance, such that the meta-learner produces $\beta$ values that are, on average, $20.11\%\pm 29.34$ lower than those of Myopic under medium spreadability (with similar results for high); we note that the wide variance in these numbers reflects the broad range of difficulty across networks in the corpus. In contrast, the fast ensemble algorithm’s performance is only $9.34\%\pm 28.34$ lower for medium spreadability (similar results for high), indicating both room for improvement by the meta-learner with a better feature set as well as an upper limit to that improvement with the current scalable algorithms. We note, however, that lower performance is not universal: for 34 and 22 networks (20% and 12.8%), the meta-learner outperforms Myopic on medium and high spreadability, respectively (Figs. 9 and S8). Under low spreadability, the fast meta-learner’s average performance generally exceeds Myopic because of the inherent precision limitation of ProbEst in this setting.
VI Discussion and Conclusion
We evaluate new and existing algorithms, and introduce a meta-learning method for choosing $k$ seed nodes to maximize the minimum access probability $\pi_{i}$ of a node in a network. The meta-learner achieves a large (75x) speedup over the existing state-of-the-art, with a modest decrease in performance over a large corpus of networks. The previous Monte Carlo method for calculating a node’s information access has high computational costs that limit the applicability of algorithms that rely on it, while the algorithms we introduced here scale-up to much larger networks.
To evaluate the performance of the algorithms, we introduce a large cross-domain network corpus, and a new performance metric $\beta$. We also introduce the concept of spreadability on a network to choose the independent cascade activation probability $\alpha$, accounting for the influence of a network’s structural features on information spreading.
Our findings have implications for the design of intervention strategies that aim to improve information access of the disadvantaged individuals in a network. This work suggests that the structure of the network plays a significant role in the performance of the algorithms, with average degree being a particularly important factor. The introduction of an ensemble method based on fast algorithms that do not rely on Monte Carlo simulation also leaves open the incorporation of future lightweight algorithms.
Acknowledgements: The authors thank Daniel B. Larremore and Zachary Kilpatrick for helpful feedback, and they acknowledge the BioFrontiers IT group at the University of Colorado Boulder for their support with data storage infrastructure, data management services, and High Performance Computing resources.
Funding: This work was supported in part by National Science Foundation Awards IIS 1956183 (DRW, CJW, AC), IIS 1956286 (ACr, BDS), and IIS 1955321 (SF).
Data and code: The data sets and code used in this study are available online at https://github.com/rhythmthief/FairnessNetworks.
Appendix A Network Corpus
Summary statistics for the network corpus can be found in Table S1. The corpus includes 174 networks drawn from 6 domains: biological (34), social (44), economic (43), technological (32), transportation (17), and information (4).
Networks were drawn from the Index of Complex Networks ICON , a large-scale index of research-quality networks spanning all domains of science. All networks included in the corpus are simple graphs, meaning their edges are undirected, unweighted, and there are no self-loops. Past work on such corpora indicates that domains (and even subdomains, e.g., online social networks vs. offline social networks) are highly distinguishable based on their structure alone ikehara2017 . Hence, a specific effort was made to (i) balance the classes, so that no domain was more than 25% of the corpus, (ii) avoid over-representing networks from particular sources (e.g., Twitter follower networks), and (iii) ensure that the minimum network size was large enough to provide good results for information spreading tasks (minimum number of nodes $n_{\min}=500$). These choices improve the breadth and variety of network structure represented in the corpus, and its utility in the analyses of this study, making a range of difficulty for the tasks.
Figure S1: Illustrations of two runs of Myopic for different initial seeds (red), with new selected seeds (yellow), and fixed $\alpha=0.5$. Numbers indicate $\pi$ for each node after the new seed is selected. Initialization significantly affects the performance of Myopic.
net. count $\langle n\rangle$ $\langle m\rangle$ $\langle k\rangle$ $\langle C\rangle$ $\langle\ell_{\rm max}\rangle$ $\langle\sigma^{2}\rangle$
full corpus 174 1495.42 26288.68 22.60 0.24 13.79 1997.12
biological 34 927.50 2830.24 7.00 0.08 11.59 285.87
social 44 1647.75 29055.57 28.11 0.42 11.02 1099.06
economic 43 2367.40 71946.23 51.97 0.38 7.19 6647.82
technological 32 843.13 1651.50 4.07 0.07 17.84 69.55
transportation 17 1243.65 2148.53 3.80 0.10 35.76 37.79
informational 4 1561.75 4124.25 6.41 0.11 8.00 174.06
Table S1: Summary statistics of the network corpus used to evaluate algorithms for the Information Access Gap Minimization problem, showing the number of networks by scientific domain, along with the average number of nodes $\langle n\rangle$, average number of edges $\langle m\rangle$, average degree $\langle k\rangle$, average clustering coefficient (transitivity) $\langle C\rangle$, average diameter $\langle\ell_{\max}\rangle$, and average variance of the degree distribution $\langle\sigma^{2}\rangle$ for networks in that domain.
Appendix B Pseudocode for Algorithms
For completeness, here we provide pseudocode for all 10 new algorithms, along with high-level descriptions. We also describe the ensemble and meta-learner approaches.
Algorithm 1: Myopic BFS
Approximates $\pi_{i}$ values by performing BFS traversal from the most recent seed and computing $\pi_{i}$ activation probabilities for every node in the network, taking into account only its “parents” and “neighbors.” For a node $x$ at distance $t$ from the seed, a “parent” is any node at distance $t-1$ from the seed that shares an edge with $x$, and a “neighbor” is any node at distance $t$ that shares an edge with $x$. When adding a new seed, update the old probabilities with a new BFS traversal.
1:$G=(V,E),s,\alpha$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   $Q\leftarrow S[-1]$ // initialize queue with latest seed
7:   while $|Q|<0$ do
8:      Perform BFS traversal and record distance $t$ to $S[-1]$
9:      for every node
10:   end while
11:   for $v\in V$ do
12:      Approximate access probability $\pi_{i}\forall v\in V$
13:      // see the writeup earlier for more info
14:   end for
15:   // Choose the node with min. activation prob.
16:   $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{Prob}(v)$
17:   $S=S\cup\{x\}$
18:end while
19:Return $S$.
List of Algorithms 1 Myopic BFS
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:$Q\leftarrow S[-1]$ // initialize queue with first random seed
6:while $|Q|<0$ do
7:   Perform BFS traversal and record distance $t$ to $S[-1]$ for every node
8:end while
9:for $v\in V$ do
10:   Approximate access probability $\pi_{i}\forall v\in V$
11:   // see the writeup earlier for more info.
12:end for
13:// Choose $s$ nodes with min. activation prob.
14:while $|S|\leq s$ do
15:   $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{Prob}(v)$
16:   $S=S\cup\{x\}$
17:end while
18:Return $S$.
List of Algorithms 2 Naive Myopic BFS
Algorithm 2: Naive Myopic BFS
Approximates $\pi_{i}$ values in the same way as Myopic BFS. Reuses the approximations from the first traversal when adding each subsequent seed.
Algorithm 3: Myopic PPR
Choose an initial seed node uniformly at random. Then, for each new seed, perform a pass of Personalized Page Rank, with personalization parameter set to bias random walk restarts from the current seed set. Sort the nodes by PPR in ascending order, choose $s$ lowest-scoring nodes as new seeds.
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{PPR}(v,S)$
7:   $S=S\cup\{x\}$
8:end while
9:Return $S$.
List of Algorithms 3 Myopic PPR
Algorithm 4: Naive Myopic PPR
Choose an initial seed node uniformly at random. Then, perform a pass of Personalized Page Rank, with personalization parameter set to bias random walk restarts from the initial random seed. Sort the nodes by PPR in ascending order, choose $s$ lowest-scoring nodes as new seeds.
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:// Sort the nodes by ascending PPR
6:$v_{1},v_{2},\ldots,v_{n}=\texttt{PPR}(G,S)$
7:$S=S\cup\{v_{1},v_{2},\ldots v_{s}\}$
8:Return $S$.
List of Algorithms 4 Naive Myopic PPR
Algorithm 5: LeastCentral
Choose an initial seed node uniformly at random. Then choose the node with the lowest closeness centrality as the new seed.
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose node with min. Close. Centrality
7:   $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{CC}(v)$
8:   $S=S\cup\{x\}$
9:end while
10:Return $S$.
List of Algorithms 5 LeastCentral
Algorithm 6: LeastCentral_n
Choose an initial seed node uniformly at random. Then choose a node with the smallest closeness centrality and select that node’s highest-degree neighbor to be the next seed.
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose node with min. Close. Centrality
7:   $x=\operatorname{argmin}_{v\in V\setminus S}\texttt{CC}(v)$
8:   // Choose the highest degree neighbor.
9:   $y=\operatorname{argmax}_{v\in N(x)}d(v)$
10:   $S=S\cup\{y\}$
11:end while
12:Return $S$.
List of Algorithms 6 LeastCentral_n
Algorithm 7: MinDegree_hc
Choose an initial seed node uniformly at random. Then, identify all minimum degree nodes, and sort by harmonic centrality. Finally, choose the node with the lowest harmonic centrality as the new seed.
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose nodes with min. degree.
7:   $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall\ w\in V\setminus S\}$
8:   // Choose node with min. Harm. Centrality.
9:   $x=\operatorname{argmin}_{v\in V^{\prime}}\texttt{HC}(v)$
10:   $S=S\cup\{x\}$
11:end while
12:Return $S$.
List of Algorithms 7 MinDegree_hc
Algorithm 8: MinDegree_hcn
Choose an initial seed node uniformly at random. Then, identify all minimum degree nodes, and sort by harmonic centrality. Finally, choose the highest-degree neighbor of the node with the lowest harmonic centrality as the new seed.
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose nodes with min. degree.
7:   $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall w\in V\setminus S\}$
8:   // Choose node with min. Harm. Centrality.
9:   $x=\operatorname{argmin}_{v\in V^{\prime}}\texttt{HC}(v)$
10:   // Choose highest degree neighbor.
11:   $y=\operatorname{argmax}_{v\in N(x)}d(v)$
12:   $S=S\cup\{y\}$
13:end while
14:Return $S$.
List of Algorithms 8 MinDegree_hcn
Algorithm 9: MinDegree_nd
Choose an initial seed node uniformly at random. Then, identify all minimum degree nodes, and sort by neighbor degree. Choose the node with the highest-degree neighbor as the new seed.
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose nodes with min. degree.
7:   $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall w\in V\setminus S\}$
8:   // Choose node with highest neigh. degree.
9:   $x=\operatorname{argmax}_{v\in V^{\prime}}\sum_{w\in N(v)}d(w)$
10:   $S=S\cup\{x\}$
11:end while
12:Return $S$.
List of Algorithms 9 MinDegree_nd
Algorithm 10: MinDegree_ndn
Choose an initial seed node uniformly at random. Then, identify all minimum degree nodes, and sort by neighbor degree. For the node with the highest-degree neighbor, choose the highest-degree neighbor itself as the new seed.
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:while $|S|\leq s$ do
6:   // Choose nodes with min. degree.
7:   $V^{\prime}=\{v\in V\setminus S\colon d(w)\geq d(v)\ \forall w\in V\setminus S\}$
8:   // Choose node with highest neigh. degree.
9:   $x=\operatorname{argmax}_{v\in V^{\prime}}\sum_{w\in N(v)}d(w)$
10:   // Choose highest degree neighbor.
11:   $y=\operatorname{argmax}_{v\in N(x)}d(v)$
12:   $S=S\cup\{y\}$
13:end while
14:Return $S$.
List of Algorithms 10 MinDegree_ndn
Algorithm 11: Fast Ensemble (Oracle)
Select five members of $D$, an ensemble of algorithms. To do so, perform greedy search for a set of five algorithms that attain $\geq 80\%$ of Myopic’s performance on the network corpus, as evaluated by ProbEst. Then, for each new network, select the algorithm that performs best as suggested by the oracle, and apply it.
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:Select $a_{1},a_{2},a_{3},a_{4},a_{5}$ to maximize number of networks in the corpus for which performance of some $a_{i}$ is $\geq 80\%$ of Myopic.
6:$D\leftarrow\{a_{1},a_{2},a_{3},a_{4},a_{5}\}$
7:// ask the oracle $O$ which algorithm to use.
8:$a_{best}\leftarrow O(G,D)$
9:// Run the algorithm to get $S$.
10:$S=S\cup a_{best}(G,s)$
11:Return $S$.
List of Algorithms 11 Fast Ensemble (Oracle)
Algorithm 12: Meta-learner
Select five members of $D$, an ensemble of algorithms. To do so, perform greedy search for a set of five algorithms that attain $\geq 80\%$ of Myopic’s performance on the network corpus, as evaluated by ProbEst. Using prior network corpus data, train a model $M$ that, given a new network $G$, selects the best algorithm out of the ensemble based on topological features. Apply the selected algorithm to $G$.
1:$G=(V,E),s$.
2:$S\subseteq V$ with $|S|=s+1$.
3:// Initialize $S$ with a random node.
4:$S=\{s_{0}\in_{R}V\}$
5:Select $a_{1},a_{2},a_{3},a_{4},a_{5}$ to maximize number of networks in the corpus for which performance of some $a_{i}$ is $\geq 80\%$ of Myopic.
6:$D\leftarrow\{a_{1},a_{2},a_{3},a_{4},a_{5}\}$
7:Train Random Forest Classifier $M$ to select the best algorithm from $D$ based on network topology, using data from the network corpus.
8:// Ask $M$ which algorithm to use.
9:$a_{best}\leftarrow M(G,D)$
10:Run the algorithm to get $S$
11:$S=S\cup a_{best}(G,s)$
12:Return $S$.
List of Algorithms 12 Meta-learner
Appendix C Supplementary Results
Algorithm performance for low and high spreadability.
Figure S2 shows the mean performance $\beta$ of all the algorithms applied to all the networks in the corpus, grouped into the six network domains: biological, social, economic, technological, transportation, and informational, for low and high spreadability settings. We observe qualitatively similar results across spreadability settings, in which Myopic is often on average the best performing algorithm. However, the relative ordering of other algorithms varies substantially across domains and spreadability settings.
Figure S2: Mean performance of the intervention algorithms on each domain in the corpus, under (a) low spreadability and (b) high spreadability settings. Medium spreadability results are given in Fig. 5.
Figure S3 expands on these results (following the experimental steps in the main text), showing for the low and high spreadability settings the performance of each algorithm relative to the performance of Myopic. Notably, in the low spreadability setting, many more algorithms fall into the “equivalent” category, in which their performance is statistically indistinguisable from Myopic ($\beta$ within one standard error). This behavior is due to the low precision of ProbEst in this setting. We find many fewer cases of equivalence in the high spreadability setting, which is qualitatively similar to the results in the medium setting, in which across networks, many alternatives perform similarly as Myopic, and in some cases outperform Myopic. These results reinforce the finding that generally no algorithm is superior to others across networks and spreadability settings.
Figure S3: Performance of intervention algorithms on the network corpus, relative to Myopic and sorted in ascending order by network size within each domain. (a): low spreadability, $12\%$ of networks have no algorithm better than or within $80\%$ of Myopic’s performance; (b): high spreadability, showing very similar results to medium spreadability (Fig. 6). “Equivalent” defined as within one standard error of $\beta$ for Myopic; typically about 0.001.
Figures S4 and S5 tabulate the counts of how often a particular algorithm was the best performing (highest $\beta$) and plot them as a function of network mean degree $\langle k\rangle$, for low and high spreadability, respectively. Here we see clearly that Myopic is by far the best algorithm in high spreadability (similar to results for medium, shown in Fig. 7), but is much less so under low spreadability. We believe this difference is attributable to the imprecision of ProbEst in the low spreadability setting.
Figure S4: Best-performing algorithm on a network vs average degree of the network under low spreadability. Each network’s corresponding single most-performant algorithm is plotted with a circle. Total number of circles per line is given on the right-hand side, indicating how many times an algorithm was the best-performing one across the entire corpus.
Figure S5: Best-performing algorithm on a network vs average degree of the network under high spreadability. Each network’s corresponding single most-performant algorithm is plotted with a circle. Total number of circles per line is given on the right-hand side, indicating how many times an algorithm was the best-performing one across the entire corpus.
Meta-learner.
To construct the meta-learner, we trained a model to predict which algorithm would perform best on a given network, given only the network’s features and the spreadability of the information in the Independent Cascade model. For this task, we held out Myopic and Naive Myopic, as the goal is to approximate their state-of-the-art performance using more scalable algorithms. Network features used are: the network domain, number of nodes $n$, average degree $\langle k\rangle$, maximum degree $k_{\max}$, degree variance $\sigma^{2}$, clustering coefficient $C$, average shortest path $\langle\ell\rangle$, diameter $\ell_{\max}$, and degree assortativity $r$. We use this approach rather than a representation learning approach to ensure better interpretability of what aspects of network structure influence performance predictability.
This multi-class classification task is challenging. Fine-tuned models come out with an accuracy of $\approx 0.51$. Figs. S6 and S7 show the learned feature importances and confusion matrix for the medium spreadability setting (similar results for low and high spreadability models). However, generally, incorrect algorithm choices by the model tend to select an alternative that is almost as good as the true best, in terms of performance $\beta$, and so classification errors often do not translate into large losses in performance.
From the classification task, we find that the domain of a network contributes the least to the prediction accuracy, followed by number of nodes $n$. Other network features are roughly equally important (Fig. S6). This results implies that network domain does not contribute much marginal information beyond a network’s features, as has been found previously by Ref. ikehara2017 . That is, network features encode much the same information as network domain.
Figure S6: Meta-learner random forest feature importance for predicting which of five algorithms is the best choice for a given network. Higher values mean a feature contributes more to the predicted performance value.
Figure S7: Confusion matrix of the best algorithm prediction model for the meta-learner.
The accuracy of the meta-learning model under medium and high spreadability (Figs. 9 and S8) are similar, but with slightly larger performance loss and substantially larger runtime speedup in the case of high, than in medium.
Unsurprisingly, the fast ensemble algorithms (which use an oracle to choose the best heuristic for a given network) in both medium and high spreadability settings outperform the meta-learner. However, the gap between fast ensemble and meta-learning is only about 10-20%, indicating that while there is room for improvement in the meta-learner’s classifier, the optimal performance is not far off. Moreover, the fast ensemble does outperform Myopic on more cases than does the meta-learning, but again, the gap is not enormous.
Figure S8: Performance difference vs. speedup for the meta-learning algorithm relative to Myopic under high spreadability, with marginal histograms. Results averaged over $1000$ meta-learner runs, and extreme outliers have been removed for visualization purposes only. Average performance loss relative to Myopic is $28.20\pm 25.80$, with an average speedup factor of $133.35\pm 79.32$, but for 22 networks (12.9%), the meta-learner outperforms Myopic.
Figure S9: Performance difference vs. speedup for the fast ensemble algorithm relative to Myopic under medium spreadability, with marginal histograms. Extreme outliers have been removed for visualization purposes only. Average performance loss relative to Myopic is $9.34\pm 28.34\%$, with an average speedup factor of $70.66\pm 85.65$, but for 48 networks (28.1%), the fast ensemble outperforms Myopic.
Figure S10: Performance difference vs. speedup for the fast ensemble algorithm relative to Myopic under high spreadability, with marginal histograms. Extreme outliers have been removed for visualization purposes only. Average performance loss relative to Myopic is $11.11\pm 21.8\%$, with an average speedup of $101.88\pm 101.27$, but for 47 networks (27.6%), the fast ensemble algorithm outperforms Myopic.